OpenStudy (anonymous):

How do you create a formula to calculate the number of people in the U.S. if 1/4th of the people equals 57.7 million?

OpenStudy (anonymous):

i have to put this in excel formula and have no clue can anyone please help

OpenStudy (anonymous):

1/4=57.7 1 = x

OpenStudy (anonymous):

im still confused im not good at math at all

OpenStudy (anonymous):

here \I got you on excel

OpenStudy (anonymous):

What do your cells look like now?

OpenStudy (anonymous):

my cells look the same

OpenStudy (anonymous):

1/4 = 57.7M 4/4 = 4 x 57.7

OpenStudy (anonymous):

Set the cell parameters as 'equation', do you know how to do that?

OpenStudy (anonymous):

http://www.twiddla.com/812747 join this i can help u guyz

OpenStudy (anonymous):

right now..its an online service u can drw and explain things

OpenStudy (anonymous):

no im totally new to all of this

OpenStudy (anonymous):

just click tht link

OpenStudy (anonymous):

Tinkerbell click that link

OpenStudy (anonymous):

http://www.twiddla.com/812747 so I can draw it for you on a spreadsheet and you can draw what you see

OpenStudy (anonymous):

im there

OpenStudy (anonymous):

Ok, all you gotta do is enter "=57.7" in one cell (lets say its G8), and then enter "=G8*4" into wherever you want the output to be

OpenStudy (anonymous):

ok so i put 57.7 in one cell then underneath i typed =(g8*4) and it gave me my answer. is that right

OpenStudy (anonymous):

Check it. 57.7 * 4 into a calculator....

OpenStudy (anonymous):

so how do u put it on the chart

OpenStudy (anonymous):

\[1/4 of x = 57\]